A poster has sides measuring 2 feet less than the sides of a square sign. If the difference between their areas is 20 square feet, find the lengths of the sides of the poster and the sign.
Answer by jorel1380(3719)

x2=(x-2)2+20
x2=x2-4x+4+20
4x=24
x=6
x-2=4
The sign is 6'x6';the poster is 4'x4'..